OpenStudy (anonymous):

If \[ a + b+ c = 0 \] then find the value of \[ {a^2 \over bc} + {b^2 \over ca} + {c^2 \over ab} ]\

OpenStudy (anonymous):

\[{a ^{2} \over bc} + {b ^{2} \over ac} + {c ^{2} \over ab} \]

OpenStudy (unklerhaukus):

\[=\frac {a^3+b^3+c^3}{abc}\]

OpenStudy (anonymous):

So the answer is 3 right?

OpenStudy (anonymous):

I got 3 too.

OpenStudy (anonymous):

ab(a+b+c) = a^2b+ab^2+abc =0 and other 2 cyclic (a+b+c)^3 = a^3 + b^3 +c^3 + 6abc + (terms equal to above including 3abc) so 3abc/abc = 3

OpenStudy (unklerhaukus):

i dont understand how ?

OpenStudy (anonymous):

If you expand (a+b+c)^3 completely you can equate all terms to 0 except 3abc

OpenStudy (unklerhaukus):

ok i follow that, but i can not see how to get (a+b+c)^3

OpenStudy (anonymous):

\[ 0=\frac{(a+b+c)^3}{a b c}=\\ \frac{a^2}{b c}+\frac{b^2}{a c}+\frac{c^2}{a b}+\frac{3 a}{b}+\frac{3 b}{a}+\frac{3 a}{c}+\frac{3 c}{a}+\frac{3 c}{b}+\frac{3 b}{c}+6\\ \frac{3 a}{b}+\frac{3 b}{a}+\frac{3 a}{c}+\frac{3 c}{a}+\frac{3 c}{b}+\frac{3 b}{c}+6=-\frac{a^2}{b c}-\frac{b^2}{a c}-\frac{c^2}{a b} \\ \frac{3 (a+b) (a+c) (b+c)}{a b c} = -\frac{a^2}{b c}-\frac{b^2}{a c}-\frac{c^2}{a b} \\ \frac{3 (-c) (-b) (-a)}{a b c} = -\frac{a^2}{b c}-\frac{b^2}{a c}-\frac{c^2}{a b} \\ -3 = -\frac{a^2}{b c}-\frac{b^2}{a c}-\frac{c^2}{a b} \\ 3 = \frac{a^2}{b c}+\frac{b^2}{a c}+\frac{c^2}{a b} \\ \]

OpenStudy (anonymous):

(a+b+c)^3 = a^3+b^3+c^3 - 3abc (after zeroing other terms per my post above) so a^3+b^3+c^3 = (a+b+c)^3 +3abc The posters original question if you add it together gives (a^3+b^3+c^3)/abc OK?

OpenStudy (unklerhaukus):

now i see , thanks eliassaab, estudier , \[\frac{(a^3+b^3+c^3)}{abc}=\frac{(a+b+c)^3 +3abc}{abc}=3\]

OpenStudy (anonymous):

OK, the numerator in the rhs comes from equating to zero all the other terms in the expansion of (a+b+c)^3...